# AN ACCELERATED PRIMAL DUAL ALGORITHM WITH BACKTRACKING FOR DECENTRALIZED CONSTRAINED OPTIMIZATION

### v1

@@
Abstract. We propose a distributed accelerated primal-dual method with backtracking (D-APDB) for cooperative multi-agent constrained consensus optimization problems over an undirected network of agents, where only those agents connected by an edge can directly communicate to exchange large-volume data vectors using a high-speed, short-range communication protocol, e.g., WiFi, and we also assume that the network allows for one-hop simple information exchange beyond immediate neighbors as in LoRaWAN protocol. The objective is to minimize the sum of agent-specific composite convex functions over agent-specific private constraint sets. Unlike existing decentralized primal-dual methods that require knowledge of the Lipschitz constants, D-APDB automatically adapts to local smoothness by employing a distributed backtracking step-size search. Each agent relies only on first-order oracles associated with its own objective and constraint functions and on local communications with the neighboring agents, without any prior knowledge of Lipschitz constants. We establish $\mathcal{O}(1/K)$ convergence guarantees for sub-optimality, infeasibility and consensus violation, under standard assumptions on smoothness and on the connectivity of the communication graph. To our knowledge, when nodes have private constraints, especially when they are nonlinear convex constraints onto which projections are not cheap to compute, D-APDB is the first distributed method with backtracking that achieves the optimal convergence rate for the class of constrained composite convex optimization problems. We also provide numerical results for D-APDB on a distributed QCQP problem illustrating the potential performance gains that can be achieved by D-APDB.
Keywords. Decentralized optimization, backtracking, parameter-free optimization methods, adaptive methods
@@
00footnotetext: The authors can be reached at the following email addresses: qjx5019@psu.edu (Q. Xu); nsa10@psu.edu (N. S. Aybat); and mg1366@rutgers.edu (M. Gürbüzbalaban).
@@ 1. Introduction
Modern datasets are large, and data are often acquired by computational agents connected over a communication network. In this context, an agent denotes any computational entity (e.g., device, node, sensor, or processor) capable of local computation and communication. In such systems, information is inherently distributed across agents that exchange messages only with neighbors, rendering centralized processing impractical or infeasible due to bandwidth, latency, energy, and privacy constraints. These factors motivate decentralized optimization algorithms, wherein data and computation reside across multiple agents and the global objective is optimized via local updates and limited message passing, without sharing raw data. Such methods arise in a wide range of applications, including machine learning with decentralized data, control and coordination in multi-robot systems, smart grids, signal processing and estimation over sensor networks [aybat2017distributed, ideal, Lu2021DecentralizedPG, dall2013distributed, rabbat2004distributed]. We refer the reader to [SYS-004, sayed2014adaptation] for additional examples in a broader context. In these settings, the goal is to collaboratively solve an optimization problem defined by agent-specific objectives and/or constraints through utilizing the computing capability of agents across the network—without a central coordinator—using only local computations and peer-to-peer communication among the neighboring nodes.
@@
where $\phi_{i}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is a proper, closed convex (possibly nonsmooth) function, $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth convex function, $g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}}$ is a smooth $\mathcal{K}_{i}$-convex function [^1] Given a closed convex cone $\mathcal{K}\subset\mathbb{R}^{m}$, and $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $\mathcal{K}$-convex if $f(\lambda x^{\prime}+(1-\lambda)x^{\prime\prime})\preceq_{\mathcal{K}}\lambda f(x^{\prime})+(1-\lambda)f(x^{\prime\prime})$ holds for all $x^{\prime},x^{\prime\prime}$ and $\lambda\in[0,1]$, where $\preceq_{\mathcal{K}}$ denotes the partial order induced by $\mathcal{K}$, i.e., $y^{\prime}\preceq_{\mathcal{K}}y^{\prime\prime}$ for $y^{\prime},y^{\prime\prime}\in\mathbb{R}^{m}$ when $y^{\prime\prime}-y^{\prime}\in\mathcal{K}$., and $\mathcal{K}_{i}\subseteq\mathbb{R}^{m_{i}}$ is a closed convex cone for $i\in\mathcal{N}$. The data of agent $i$ defining $\varphi_{i}$, $g_{i}$ and $\mathcal{K}_{i}$ are private and not shared globally, and agents cooperate to solve (1.1) using only local computation and limited communication. This is a general class of problems that include many applications and important special cases [Lu2021DecentralizedPG, Nedi2009DistributedSM, Boyd2011DistributedOA, lee-nedich, SYS-004, sayed2014adaptation].
@@
Over the past few decades, a rich literature has emerged on decentralized constrained optimization problems that are special cases of the class of problems in (1.1). In particular, distributed (sub)gradient schemes, gradient tracking methods, and their accelerated and proximal variants have been extensively studied under a variety of assumptions on smoothness, convexity, and network connectivity, e.g., [ghaderyan2023fast, 5404774, 7169615, 8752033, NEURIPS2020_d4b5b5c1, Jakoveti2011FastDG, shi2015extra]. Other approaches include primal–dual type methods based on Augmented Lagrangian and ADMM formulations, e.g., [aybat2016primal, aybat2017distributed, aybat2015asynchronous, ghaderyan2023fast, koshal2011multiuser, ideal, hamedani2021decentralized, Boyd2011DistributedOA]. The latter class of algorithms typically rely on Lagrangian formulations and exploit separability across agents to achieve fully distributed implementations. Except very few [aybat2016primal, hamedani2021decentralized], none of the aforementioned methods can handle (1.1) in its full generality without employing projections onto $\{x:-g_{i}(x)\in\mathcal{K}_{i}\}$. More importantly, a persistent challenge in nearly all of these methods is to appropriately choose stepsizes with theoretical convergence guarantees. Such guarantees often require stepsizes that depend on global problem parameters—such as global Lipschitz constants for $\nabla f_{i}$ and Jacobians of $g_{i}$ for $i\in\mathcal{N}$, that are difficult to estimate, rarely known a-priori, and may vary widely across agents. In practice, this leads practitioners to grid-search or adopt conservative, globally synchronized stepsizes that can significantly degrade the convergence speed. The difficulty is further exacerbated in the presence of nonlinear functional constraints, where the relevant Lipschitz constants for the Lagrangian gradients depend on the dual iterates, which typically evolve within a cone; therefore, the Lipschitz constants may be only locally finite but not globally.
@@
The literature on parameter-free decentralized algorithms—i.e., methods that do not require prior knowledge of Lipschitz constants—remains limited. A small body of work studies decentralized adaptive gradient methods that exploit past gradient information to choose stepsizes adaptively across iterations. For example, [nazari2022dadam] develops a distributed adaptive moment estimation method for online smooth convex and nonconvex minimization and establishes regret bounds, while [chen2023decentralizedadaptive] proposes a general framework for converting centralized adaptive gradient schemes such as RMSProp, Adam, and AdaGrad into decentralized counterparts. We should emphasize that for the methods proposed in [chen2023decentralizedadaptive] to have convergence guarantees, it is required that the step size is sufficiently small inversely proportional to $1/\max_{i=1,\ldots,N}\{L_{f_{i}}\}$, where $L_{f_{i}}$ is the Lipschitz constant of $\nabla f_{i}$, and this would require some information exchange beyond the immediate neighbors on the network. Furthermore, both [nazari2022dadam] and [chen2023decentralizedadaptive] consider unconstrained smooth optimization problems and do not handle constraints.
@@
The work [li2024problem] proposes a parameter-free decentralized algorithm for nonconvex stochastic optimization that does not require knowledge of $\{L_{f_{i}}\}_{i\in\mathcal{N}}$ or topological information about the communication network. Another line of work [aldana2024towards] introduces a port-Hamiltonian systems framework for the design and analysis of distributed optimization algorithms, and develops methods for smooth, strongly convex, and unconstrained decentralized problems. Convergence of the proposed decentralized methods is guaranteed for certain special graph classes or when the stepsize is chosen below a threshold that depends on global network information [aldana2024towards, Corollaries 2 and 3]. However, we should emphasize that the proposed approaches in [li2024problem, aldana2024towards] both focus on smooth minimization problems and they do not address either the node-specific constraints or the non-smooth terms in the objective.
@@
More recently, [kuruzov2024achieving] considered distributed smooth strongly convex unconstrained minimization problems and proposed a parameter-free decentralized optimization algorithm that employs a local backtracking line search to select stepsizes adaptively—without requiring global information or heavy communication, and the authors analyze two different implementations of the proposed method. The first implementation uses a global min-consensus to synchronize agents’ stepsizes, which can be computed in practice via flooding protocols over low-power wide-area networks, e.g., LoRa [askhedkar2020hardware, janssen2020lora, kim2016low] (low power, long range), as it requires simple information exchange beyond immediate neighbors. On the other hand, the second one replaces the global min-consensus update with a local min-consensus update, restricting communication to immediate neighbors; however, as mentioned in the follow-up work [kuruzov2025adaptive], this increased locality comes at the cost of weaker theoretical guarantees due to stronger assumptions requiring that the iterates stay bounded, and potentially non-monotone convergence trajectories. In [kuruzov2025adaptive], the authors introduces a new fully decentralized algorithm in which each agent adaptively selects its own stepsize using only neighbor-to-neighbor communication and no global information is required for its implementation—agents need not even know whether the problem is strongly convex. The algorithm retains strong guarantees: it converges at a linear rate under strong convexity and at a $\mathcal{O}(1/T)$-sublinear rate under mere convexity, matching the best-known rates for parameter-dependent, nonadaptive distributed methods. In the merely convex setting, the convergence guarantees for the method in [kuruzov2025adaptive] also require that the primal-dual iterate sequence stays bounded (see [kuruzov2025adaptive, Theorem 10]) as in their earlier work [kuruzov2024achieving] –since neither work [kuruzov2024achieving, kuruzov2025adaptive] can handle indicator functions to directly impose boundedness, the authors argue that this requirement can be satisfied by some particular choice of algorithmic parameter sequences.
@@
All of the previously mentioned works are designed to address unconstrained decentralized problems of the form $\min_{x\in\mathbb{R}^{n}}\sum_{i\in\mathcal{N}}f_{i}(x)$, where each $f_{i}$ is smooth with a Lipschitz-continuous gradient. Consequently, they do not apply to the more general constrained problems we consider in (1.1), which feature agent-specific local constraints and nonsmooth terms $\phi_{i}(x)$ in the objective. A more recent work [chen2025parameter] is the first to accommodate nonsmoothness, focusing on problems of the form $\min\sum_{i\in\mathcal{N}}f_{i}(x)+\phi(x)$, where $\phi$ is a convex, nonsmooth function known to all agents in the network. The work in [chen2025parameter] introduces a novel three-operator splitting technique and proposes a method that require neither global network information nor extensive inter-agent communication. The resulting adaptive decentralized method enjoys robust convergence guarantees and outperforms existing nonadaptive approaches.
@@
These aforementioned results on adaptive distributed methods are exciting developments for decentralized optimization; nevertheless, none of these existing methods that we discussed above can handle a problem of the form (1.1) in its full generality. Indeed, to our knowledge, there is no decentralized optimization method that does not rely on the a-priori knowledge of Lipschitz constants and that can handle either node-specific nonsmooth closed convex terms in the objective function or agent-specific local constraints that are defined by nonlinear convex functions.
@@
We should mention that for centralized unconstrained optimization, where a central node can aggregate and process all data, a wide range of parameter-free methods is available. These include centralized adaptive gradient schemes such as AdaGrad [duchi2011adagrad], RMSProp [tieleman2012rmsprop], Adam [kingma2015adam], and their variants [reddi2018convergence, luo2019adabound], as well as methods based on Barzilai–Borwein stepsizes [zhou2025adabb, barzilai1988two, burdakov2019stabilized] and adaptive techniques that estimate local curvature [malitsky2024adaptive, malitsky2020adaptive]. For centralized constrained convex optimization, backtracking schemes can estimate local Lipschitz constants on the fly, yielding convergence-rate guarantees without prior knowledge of problem-specific smoothness parameters. In particular, Lagrangian reformulations of (1.1) lead to convex–concave saddle–point (SP) problems, for which primal–dual methods with backtracking have been developed [hamedani2021primal, jiang_mokthari] –on a different note, there are also backtracking-based algorithms for centralized non-convex min–max formulations [xu2024stochastic, zhang2024agda+], which can be used for (1.1) as well. However, these methods fundamentally rely on centralized aggregation; therefore, they do not extend directly to the decentralized computation setting we consider in this paper, where no single node has access to global information. There are also adaptive distributed methods in federated learning or central-server settings, e.g., [chen2020toward, Xie2020, reddi2021adaptive]; but these algorithms still require a central node to aggregate information from the network and are thus not applicable to the fully decentralized setting over an arbitrary connected undirected communication network considered in this paper.
@@
To address this gap, we propose D-APDB, a novel decentralized primal–dual algorithm that incorporates a backtracking mechanism for local, agent-specific stepsize selection to solve (1.1). Unlike existing decentralized primal-dual methods that require knowledge of the Lipschitz constants, D-APDB automatically adapts to local smoothness by employing a distributed backtracking step-size search. Each agent relies only on first-order oracles associated with its own objective and constraint functions, without any prior knowledge of Lipschitz constants. At each iteration of D-APDB, each agent locally communicates a $n$-dimensional vector one time with the neighboring agents, and the agents across the network collectively implement a max-consensus one time. As explained in [kuruzov2024achieving], max-consensus protocol is well-suited to existing wireless mesh network technologies; more precisely, LoRa [askhedkar2020hardware, janssen2020lora, kim2016low] enables wide-area coverage at low data rates, which is advantageous for network-wide flooding where each transmission reaches all nodes in one hop but conveys limited information. We establish $\mathcal{O}(1/K)$ convergence guarantees for sub-optimality, infeasibility and consensus violation, under standard assumptions on smoothness and on the connectivity of the communication graph. To our knowledge, D-APDB is the first distributed method with backtracking that achieves the optimal convergence rate for the class of composite convex optimization problems subject to functional agent-specific convex constraints. Furthermore, we propose a variant of our method, which we call D-APDB0, tailored to the setting with $g_{i}(\cdot)=0$ for all $i\in\mathcal{N}$ while agents can still have closed convex functions $\phi_{i}$ in their local objectives. D-APDB0 can also achieve $\mathcal{O}(1/K)$ convergence guarantees for sub-optimality and consensus violation –to the best our knowledge, this is the first time a rate result is provided for a decentralized method that can handle node-specific closed convex functions without requiring apriori knowledge on Lipschitz constants. Finally, we present numerical results for D-APDB0 and D-APDB on a distributed $\ell_{1}$-norm regularized QP and QCQP problems, illustrating the potential performance gains achievable with our proposed backtracking framework.
@@
Throughout the text $\|\cdot\|$ denotes the Euclidean norm and the spectral norm when the argument is a vector and a matrix, respectively. Given two sets $A$ and $B$, $A\times B$ denotes the Cartesian product, and $\Pi_{i=1}^{M}A_{i}$ denotes the Cartesian product of a collection of sets $\{A_{i}\}_{i=i}^{M}$. In few places, we abuse the notation $\Pi_{i=1}^{M}a_{i}$ to denote the multiplication of a collection of real numbers $\{a_{i}\}_{i=1}^{M}\subset\mathbb{R}$. Given a set $S$, let $\text{1}_{S}(\cdot)$ denotes its indicator function, i.e., $\text{1}_{S}(x)=0$ if $x\in S$ and is equal to $\infty$ otherwise. For $S\subset\mathbb{R}^{n}$, $\mathcal{P}_{S}(\cdot)$ denoted the Euclidean projection onto $S$; moreover, $d_{S}:\mathbb{R}^{n}\to\mathbb{R}_{+}$ denotes the distance function, i.e., $d_{S}(x)=\|x-\mathcal{P}_{S}(x)\|$. Given a convex cone $\mathcal{K}\subseteq\mathbb{R}^{m}$, let $\mathcal{K}^{*}$ denote its dual cone, i.e., $\mathcal{K}^{*}\triangleq\{\theta\in\mathbb{R}^{m}:\langle\theta,w\rangle\geq 0,\quad\forall w\in\mathcal{K}\}$. Given a discrete set $\mathcal{S}$, $|\mathcal{S}|$ denotes the cardinality of $\mathcal{S}$. The set $\mathbb{S}^{n_{x}}_{++}$ denotes the set of $n_{x}\times n_{x}$ positive definite matrices. Throughout the text, we use $\mathop{\bf diag}(\cdot)$ to construct block-diagonal matrices, i.e., given $G_{i}\in\mathbb{R}^{m_{i}\times n}$ for $i\in\mathcal{N}$, $G=\mathop{\bf diag}\left([G_{i}]_{i\in\mathcal{N}}\right)$ denotes a block diagonal matrix with diagonal blocks being $G_{i}$ for $i\in\mathcal{N}$; hence, $G\in\mathbb{R}^{m\times n|\mathcal{N}|}$ where $m=\sum_{i\in\mathcal{N}}m_{i}$. We use $I_{m}$ to denote the $m\times m$ identity matrix, and $\mathbf{1}_{m}$ to denote the $m$-dimensional vector of ones.

### v2

@@
Abstract. We propose a distributed accelerated primal-dual method with backtracking (D-APDB) for cooperative multi-agent constrained consensus optimization problems over an undirected network of agents, where only those agents connected by an edge can directly communicate to exchange large-volume data vectors using a high-speed, short-range communication protocol, e.g., WiFi, and we also assume that the network allows for one-hop simple information exchange beyond immediate neighbors as in LoRaWAN protocol. The objective is to minimize the sum of agent-specific composite convex functions over agent-specific private constraint sets. Unlike existing decentralized primal-dual methods that require knowledge of the Lipschitz constants, D-APDB automatically adapts to local smoothness by employing a distributed backtracking step-size search. Each agent relies only on first-order oracles associated with its own objective and constraint functions and on local communications with the neighboring agents, without any prior knowledge of Lipschitz constants. We establish $\mathcal{O}(1/K)$ convergence guarantees for sub-optimality, infeasibility and consensus violation, under standard assumptions on smoothness and on the connectivity of the communication graph. To our knowledge, when nodes have private constraints, especially when they are nonlinear convex constraints onto which projections are not cheap to compute, D-APDB is the first distributed method with backtracking that achieves the optimal convergence rate for the class of constrained composite convex optimization problems. We provide numerical results for D-APDB on a distributed QCQP problem and distributed primal SVM training, illustrating the potential performance gains that can be achieved by D-APDB.
Keywords. Adaptive methods; Backtracking; Decentralized optimization; Parameter-free optimization methods.
@@
footnotetext: Corresponding author. E-mail addresses: qjx5019@psu.edu (Q. Xu), nsa10@psu.edu (N.S. Aybat), mg1366@rutgers.edu (M. Gürbüzbalaban).
@@ 1. Introduction
Modern datasets are large, and data are often acquired by computational agents connected over a communication network. In this context, an agent denotes any computational entity (e.g., device, node, sensor, or processor) capable of local computation and communication. In such systems, information is inherently distributed across agents that exchange messages only with neighbors, rendering centralized processing impractical or infeasible due to bandwidth, latency, energy, and privacy constraints. These factors motivate decentralized optimization algorithms, wherein data and computation reside across multiple agents and the global objective is optimized via local updates and limited message passing, without sharing raw data. Such methods arise in a wide range of applications, including machine learning with decentralized data, control and coordination in multi-robot systems, smart grids, signal processing and estimation over sensor networks [2, 7, 15, 33, 42]. We refer the reader to [40, 46] for additional examples in a broader context. In these settings, the goal is to collaboratively solve an optimization problem defined by agent-specific objectives and/or constraints through utilizing the computing capability of agents across the network—without a central coordinator—using only local computations and peer-to-peer communication among the neighboring nodes.
@@
where $\phi_{i}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is a proper, closed convex (possibly nonsmooth) function, $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth convex function, $g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{m_{i}}$ is a smooth $\mathcal{K}_{i}$-convex function11Given a closed convex cone $\mathcal{K}\subset\mathbb{R}^{m}$, and $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is $\mathcal{K}$-convex if $f(\lambda x^{\prime}+(1-\lambda)x^{\prime\prime})\preceq_{\mathcal{K}}\lambda f(x^{\prime})+(1-\lambda)f(x^{\prime\prime})$ holds for all $x^{\prime},x^{\prime\prime}$ and $\lambda\in[0,1]$, where $\preceq_{\mathcal{K}}$ denotes the partial order induced by $\mathcal{K}$, i.e., $y^{\prime}\preceq_{\mathcal{K}}y^{\prime\prime}$ for $y^{\prime},y^{\prime\prime}\in\mathbb{R}^{m}$ when $y^{\prime\prime}-y^{\prime}\in\mathcal{K}$., and $\mathcal{K}_{i}\subseteq\mathbb{R}^{m_{i}}$ is a closed convex cone for $i\in\mathcal{N}$. The data of agent $i$ defining $\varphi_{i}$, $g_{i}$ and $\mathcal{K}_{i}$ are private and not shared globally, and agents cooperate to solve (1.1) using only local computation and limited communication. This is a general class of problems that include many applications and important special cases [9, 30, 33, 39, 40, 46].
@@
Over the past few decades, a rich literature has emerged on decentralized constrained optimization problems that are special cases of the class of problems in (1.1). In particular, distributed (sub)gradient schemes, gradient tracking methods, and their accelerated and proximal variants have been extensively studied under a variety of assumptions on smoothness, convexity, and network connectivity, e.g., [18, 22, 32, 38, 49, 50, 54]. Other approaches include primal–dual type methods based on Augmented Lagrangian and ADMM formulations, e.g., [2, 4, 5, 7, 9, 18, 19, 27]. The latter class of algorithms typically rely on Lagrangian formulations and exploit separability across agents to achieve fully distributed implementations. Except very few [5, 19], none of the aforementioned methods can handle (1.1) in its full generality without employing projections onto $\{x:-g_{i}(x)\in\mathcal{K}_{i}\}$. More importantly, a persistent challenge in nearly all of these methods is to appropriately choose stepsizes with theoretical convergence guarantees. Such guarantees often require stepsizes that depend on global problem parameters—such as global Lipschitz constants for $\nabla f_{i}$ and Jacobians of $g_{i}$ for $i\in\mathcal{N}$, that are difficult to estimate, rarely known a-priori, and may vary widely across agents. In practice, this leads practitioners to grid-search or adopt conservative, globally synchronized stepsizes that can significantly degrade the convergence speed. The difficulty is further exacerbated in the presence of nonlinear functional constraints, where the relevant Lipschitz constants for the Lagrangian gradients depend on the dual iterates, which typically evolve within a cone; therefore, the Lipschitz constants may be only locally finite but not globally.
@@
The literature on parameter-free decentralized algorithms—i.e., methods that do not require prior knowledge of Lipschitz constants—remains limited. A small body of work studies decentralized adaptive gradient methods that exploit past gradient information to choose stepsizes adaptively across iterations. For example, [37] develops a distributed adaptive moment estimation method for online smooth convex and nonconvex minimization and establishes regret bounds, while [12] proposes a general framework for converting centralized adaptive gradient schemes such as RMSProp, Adam, and AdaGrad into decentralized counterparts. We should emphasize that for the methods proposed in [12] to have convergence guarantees, it is required that the step size is sufficiently small inversely proportional to $1/\max_{i=1,\ldots,N}\{L_{f_{i}}\}$, where $L_{f_{i}}$ is the Lipschitz constant of $\nabla f_{i}$, and this would require some information exchange beyond the immediate neighbors on the network. Furthermore, both [37] and [12] consider unconstrained smooth optimization problems and do not handle constraints.
@@
The work [31] proposes a parameter-free decentralized algorithm for nonconvex stochastic optimization and the proposed method does not require knowledge of $\{L_{f_{i}}\}_{i\in\mathcal{N}}$ or topological information about the communication network. Another line of work [1] introduces a port-Hamiltonian systems framework for the design and analysis of distributed optimization algorithms, and develops methods for smooth, strongly convex, and unconstrained decentralized problems. Convergence of the proposed decentralized methods is guaranteed for certain special graph classes or when the stepsize is chosen below a threshold that depends on global network information [1, Corollaries 2 and 3]. However, we should emphasize that the proposed approaches in [1, 31] both focus on smooth minimization problems and they do not address either the node-specific constraints or the non-smooth terms in the objective.
@@
More recently, [29] considered distributed smooth strongly convex unconstrained minimization problems and proposed a parameter-free decentralized optimization algorithm that employs a local backtracking line search to select stepsizes adaptively—without requiring global information or heavy communication, and the authors analyze two different implementations of the proposed method. The first implementation uses a global min-consensus to synchronize agents’ stepsizes, which can be computed in practice via flooding protocols over low-power wide-area networks, e.g., LoRa [3, 23, 25] (low power, long range), as it requires simple information exchange beyond immediate neighbors. On the other hand, the second one replaces the global min-consensus update with a local min-consensus update, restricting communication to immediate neighbors; however, as mentioned in the follow-up work [28], this increased locality comes at the cost of weaker theoretical guarantees due to stronger assumptions requiring that the iterates stay bounded, and potentially non-monotone convergence trajectories. In [28], the authors introduce a fully decentralized algorithm in which each agent adaptively selects its own stepsize using only neighbor-to-neighbor communication and no global information is required for its implementation—agents need not even know whether the problem is strongly convex. The algorithm retains strong guarantees: it converges at a linear rate under strong convexity and at a $\mathcal{O}(1/T)$-sublinear rate under mere convexity, matching the best-known rates for parameter-dependent, nonadaptive distributed methods. In the merely convex setting, the convergence guarantees for the method in [28] also require that the primal-dual iterate sequence stays bounded (see [28, Theorem 10]) as in their earlier work [29] –since neither work [28, 29] can handle indicator functions to directly impose boundedness, the authors argue that this requirement can be satisfied by some particular choice of algorithmic parameter sequences.
@@
All of the previously mentioned works are designed to address unconstrained decentralized problems of the form $\min_{x\in\mathbb{R}^{n}}\sum_{i\in\mathcal{N}}f_{i}(x)$, where each $f_{i}$ is smooth with a Lipschitz-continuous gradient. Consequently, they do not apply to the more general constrained problems we consider in (1.1), which feature agent-specific local constraints and nonsmooth terms $\phi_{i}(x)$ in the objective. A more recent work [14] is the first to accommodate nonsmoothness, focusing on problems of the form $\min\sum_{i\in\mathcal{N}}f_{i}(x)+\phi(x)$, where $\phi$ is a convex, nonsmooth function known to all agents in the network. The work in [14] introduces a novel three-operator splitting technique and proposes a method that requires neither global network information nor extensive inter-agent communication. The resulting adaptive decentralized method enjoys robust convergence guarantees and outperforms existing nonadaptive approaches.
@@
These aforementioned results on adaptive distributed methods are exciting developments for decentralized optimization; nevertheless, none of these existing methods that we discussed above can handle a problem of the form (1.1) in its full generality. Indeed, to our knowledge, there is no decentralized optimization method that does not rely on the a priori knowledge of Lipschitz constants and that can handle agent-specific local constraints defined by nonlinear convex functions.
@@
We should mention that for centralized unconstrained optimization, where a central node can aggregate and process all data, a wide range of parameter-free methods is available. These include centralized adaptive gradient schemes such as AdaGrad [16], RMSProp [51], Adam [26], and their variants [34, 44], as well as methods based on Barzilai–Borwein stepsizes [8, 10, 57] and adaptive techniques that estimate local curvature [35, 36]. For centralized constrained convex optimization, backtracking schemes can estimate local Lipschitz constants on the fly, yielding convergence-rate guarantees without prior knowledge of problem-specific smoothness parameters. In particular, Lagrangian reformulations of (1.1) lead to convex–concave saddle–point (SP) problems, for which primal–dual methods with backtracking have been developed [20, 24] –on a different note, there are also backtracking-based algorithms for centralized non-convex min–max formulations [53, 55], which can be used for (1.1) as well. However, these methods fundamentally rely on centralized aggregation; therefore, they do not extend directly to the decentralized computation setting we consider in this paper, where no single node has access to global information. There are also adaptive distributed methods in federated learning or central-server settings, e.g., [13, 43, 52]; but these algorithms still require a central node to aggregate information from the network and are thus not applicable to the fully decentralized setting over an arbitrary connected undirected communication network considered in this paper.
@@
To address this gap, we propose D-APDB, a novel decentralized primal–dual algorithm that incorporates a backtracking mechanism for local, agent-specific stepsize selection to solve (1.1). Unlike existing decentralized primal-dual methods that require knowledge of the Lipschitz constants, D-APDB automatically adapts to local smoothness by employing a distributed backtracking step-size search. Each agent relies only on first-order oracles associated with its own objective and constraint functions, without any prior knowledge of Lipschitz constants. At each iteration of D-APDB, each agent locally communicates a $n$-dimensional vector one time with the neighboring agents, and the agents across the network collectively implement a max-consensus one time. As explained in [29], the max-consensus protocol is well-suited to existing wireless mesh network technologies; more precisely, LoRa [3, 23, 25] enables wide-area coverage at low data rates, which is advantageous for network-wide flooding where each transmission reaches all nodes in one hop but conveys limited information. We establish $\mathcal{O}(1/K)$ convergence guarantees for sub-optimality, infeasibility and consensus violation, under standard assumptions on smoothness and on the connectivity of the communication graph. To our knowledge, D-APDB is the first distributed method with backtracking that achieves the optimal convergence rate for the class of composite convex optimization problems subject to functional agent-specific convex constraints. Furthermore, we propose a variant of our method, which we call D-APDB0, tailored to the setting with $g_{i}(\cdot)=0$ for all $i\in\mathcal{N}$ while agents can still have closed convex functions $\phi_{i}$ in their local objectives. D-APDB0 can also achieve $\mathcal{O}(1/K)$ convergence guarantees for sub-optimality and consensus violation –to the best of our knowledge, this is the first time a rate result is provided for a decentralized method that can handle node-specific closed convex functions without requiring a priori knowledge on Lipschitz constants. Finally, we present numerical results for D-APDB0 and D-APDB on a distributed $\ell_{1}$-norm regularized QP and QCQP problems, illustrating the potential performance gains achievable with our proposed backtracking framework.
@@
Throughout the text $\|\cdot\|$ denotes the Euclidean norm and the spectral norm when the argument is a vector and a matrix, respectively. Given two sets $A$ and $B$, $A\times B$ denotes the Cartesian product, and $\Pi_{i=1}^{M}A_{i}$ denotes the Cartesian product of a collection of sets $\{A_{i}\}_{i=i}^{M}$. In few places, we abuse the notation $\Pi_{i=1}^{M}a_{i}$ to denote the multiplication of a collection of real numbers $\{a_{i}\}_{i=1}^{M}\subset\mathbb{R}$. Given a set $S$, let $\mathds{1}_{S}(\cdot)$ denote its indicator function, i.e., $\mathds{1}_{S}(x)=0$ if $x\in S$ and is equal to $\infty$ otherwise. For $S\subset\mathbb{R}^{n}$, $\mathcal{P}_{S}(\cdot)$ denotes the Euclidean projection onto $S$; moreover, $d_{S}:\mathbb{R}^{n}\to\mathbb{R}_{+}$ denotes the distance function, i.e., $d_{S}(x)=\|x-\mathcal{P}_{S}(x)\|$. Given a convex cone $\mathcal{K}\subseteq\mathbb{R}^{m}$, let $\mathcal{K}^{*}$ denote its dual cone, i.e., $\mathcal{K}^{*}\triangleq\{\theta\in\mathbb{R}^{m}:\langle\theta,w\rangle\geq 0,\quad\forall w\in\mathcal{K}\}$. Given a discrete set $\mathcal{S}$, $|\mathcal{S}|$ denotes the cardinality of $\mathcal{S}$. The set $\mathbb{S}^{n_{x}}_{++}$ denotes the set of $n_{x}\times n_{x}$ positive definite matrices. Throughout the text, we use $\mathop{\bf diag}(\cdot)$ to construct block-diagonal matrices, i.e., given $G_{i}\in\mathbb{R}^{m_{i}\times n}$ for $i\in\mathcal{N}$, $G=\mathop{\bf diag}\left([G_{i}]_{i\in\mathcal{N}}\right)$ denotes a block diagonal matrix with diagonal blocks being $G_{i}$ for $i\in\mathcal{N}$; hence, $G\in\mathbb{R}^{m\times n|\mathcal{N}|}$ where $m=\sum_{i\in\mathcal{N}}m_{i}$. We use $I_{m}$ to denote the $m\times m$ identity matrix, and $\mathbf{1}_{m}$ to denote the $m$-dimensional vector of ones.